A rectangular garden is twice as long as it is wide and is surrounded by a sidewalk that is 4 ft wide. The area of the sidewalk is 1024 ft sq. Find the
dimensions of the garden.

Answer by mananth(16946)

width of garden x
length of garden= 2 x
sidewalk width= 4
Area of garden + walkaway= x*( 2 x)
Area of sidewalk 1024 M^2
with including walkway x + 8
length including walkway(2x +8 )
Total area – Area ofgarden = Area of sidewalk

Total area = (x+8)*( 2x+8)

(x+8)*( 2x+8)- x*(2x)=1024

2x^2+8x+16x+64-2x^2=1024

x^2 + 24 x + 64 = 1024

24 x = 960
/ 24
x= 40 Meters width of garden
80 Meters length of garden
CHECK

48 * 88 = 4224 M^2
40 * 80 = 3200 M^2
1024 M^2